What is a partially ordered set? Prove that the inclusion relation `is a subset of’ is a partial ordering on the power set of a finite set S.
1
Expert's answer
2020-06-08T19:25:17-0400

A partially ordered set is a set "X"  with a partial ordering "\leq"  on "X" . A partial ordering "\leq"  on "X"  is a binary relation on "X"  satisfying the following properties:

  • for all "x, y, z\in X" , if "x\leq y"  and "y\leq z" , then "x\leq z"  (transitivity)
  • for all "x\in X" , "x\leq x"  (reflexivity)
  • for all "x, y\in X" , if "x\leq y"  and "y\leq x" , then "x=y"  (antisymmetry)


Let "\mathcal{P}(S)"  denote the power set of "S" . Below we show that "\subseteq"  on "\mathcal{P}(S)"  is a partial ordering.

  • Transitivity. Let "T_0, T_1, T_2\in \mathcal{P}(S)" . Assume that "T_0\subseteq T_1"  and "T_1\subseteq T_2" . For all "t\in T_0" , "t\in T_1"  because "T_0\subseteq T_1" , and "t\in T_2"  because "T_1\subseteq T_2" . Therefore, "T_0\subseteq T_2" .
  • Reflexivity. Let "T\in \mathcal{P}(S)" . For all "t\in T" , "t\in T"  is true. Hence "T\subseteq T" .
  • Antisymmetry. Let "T_0, T_1\in \mathcal{P}(S)" . Assume that "T_0\subseteq T_1"  and "T_1\subseteq T_0" . Then "T_0=T_1"  by the definition of set equality.